Perform the indicated operation and simplify.

[(x+1)/(x^2-6x+8)] / [(3)/(x^2-16)]
Answer #1

Given (4+1) (x² - 6x+8) 3(x²-16) = (2+1) (x² - 2a-4x+8) 3(x2-4) = (x+1) (x(x-2) - 4(x-2)) 342+4) (2-4) = (x+1) (1-2)(x-4) 3(x